# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ Appendix B An extended discussion on the choice of the metric
B.1 Linear problems
Setting B.1.
@@ B.1 Linear problems
Choosing on $\mathcal{M}=\mathrm{V}$ the (point-independent) metric
@@
Remark B.3.
In Setting B.1, it is typically straightforward to approximate numerically the inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}}$ and the duality pairing $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}$, but not the dual inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}^{\prime}}$. Let $\mathcal{R}_{\mathrm{W}}:\mathrm{W}\to\mathrm{W}^{\prime}$ denote the Riesz map, defined by
@@ B.2 Nonlinear problems
Setting B.4.
Let $(\mathcal{M},a)$ be a Riemannian manifold and $f:\mathcal{M}\to\mathbb{R}$ a smooth function. We aim to solve $\inf_{u\in\mathcal{M}}E(u)$.
If $\mathrm{Hess}^{a}E(u)$ is symmetric positive definite for all $u\in\mathcal{M}$, a natural metric leading to Riemannian Newton’s method [14] is
@@
In the case $(\mathcal{M},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this yields the Energy Natural Gradient Descent (ENGD) method [54, 55]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathcal{M}$, as otherwise (19) does not define a valid metric.
@@
Appendix C Proof of 3.2
See 3.1
See 3.2
Proof.
Let $\mathbf{e}_{h}\in\mathbb{R}^{p}$ denote the $h$-th vector of the canonical basis, i.e., the vector that has a $1$ in the $h$-th entry and 00 in all other entries. Using the linearity and continuity of $\mathcal{F}_{u_{\boldsymbol{\theta}}}$,
or, in matrix form,
where the Hilbert adjoint $\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}$ is taken with respect to the inner product on $\mathrm{L}^{2}(\Omega,\varphi)^{k}$, and we identify a matrix with the induced linear map.
and thus, in matrix form:
Remark C.1.
More in general, it can be shown that if the metric $g_{u}$ takes the form
where $\mathcal{F}_{u}^{i}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi)^{k}$ are linear and continuous for $i=1,2$, and the integral is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}_{1}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}_{2}(\boldsymbol{\theta}),\quad\,\,\,\mathbf{F}_{i}(\boldsymbol{\theta})=\begin{bmatrix}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{i}u_{\boldsymbol{\theta}}(\mathbf{x}_{1})\\ \vdots\\ \mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{i}u_{\boldsymbol{\theta}}(\mathbf{x}_{q})\end{bmatrix}.\end{split}$ (23)
Appendix D Modification of jaxopt.BFGS

### v3

@@ Appendix A Choice of the Metric
A.1 Linear problems
Setting A.1.
@@ A.1 Linear problems
Choosing on $\mathrm{V}=\mathrm{V}$ the (point-independent) metric
@@
Remark A.3.
In Setting A.1, it is typically straightforward to approximate numerically the inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}}$ and the duality pairing $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}$, but not the dual inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}^{\prime}}$. Let $\mathcal{R}_{\mathrm{W}}:\mathrm{W}\to\mathrm{W}^{\prime}$ denote the Riesz map, defined by
@@ A.2 Nonlinear problems
Setting A.4.
Let $(\mathrm{V},a)$ be a Riemannian manifold and $f:\mathrm{V}\to\mathbb{R}$ a smooth function. We aim to solve $\inf_{u\in\mathrm{V}}E(u)$.
If $\mathrm{Hess}^{a}E(u)$ is symmetric positive definite for all $u\in\mathrm{V}$, a natural metric leading to Riemannian Newton’s method [15] is
@@
In the case $(\mathrm{V},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this yields the Energy Natural Gradient Descent (ENGD) method [64, 65]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathrm{V}$, as otherwise (31) does not define a valid metric.
@@
Appendix B Modification of jaxopt.BFGS